Test: Exam#2 Submit Test sti This Question: 1 pt 10 of 19 (13 complete) This Test: 19 pls possible n Help In a survey of 3492 adults, 1453 say they have started paying bills online in the last year Construct a 99% confidence interval for the population proportion interpret the results. A 99% confidence interval for the population proportion is Round to three decimal places as needed.) Interpret your results. Choose the correct answer below 0 A with 99% confidence it can be said that the sample proportion of adults who say they have started paying bills online in the last year s between the endpoints of the given confidence interval With 99% confidence it can be said that the population proportion of adults who say they have started paying bills online in the last year is between the endpoints of the given confidence interval B C. The endpoints of the given confidence interval show that adults pay bills online 99% of the time. Click to select your answer(s)
Answer #1

This is a sample proportion problem which can solved easily using z table and finding margin of error.

Given data,

Sample size, n = 3492

People wo have started paying bills online, x = 1453

So, the sample proportion, widehat{p} = rac{x}{n} = 1453 3492 = 0.416

The confidence interval(C.I.) for population proportion is given by-

C.I. = widehat{p} pm zsqrt{rac{widehat{p}(1-widehat{p})}{n}} where z is the z score .

For 99% confidence interval, z = 2.58

Now, lPI 1 -P 7l = 0.416(1- 0.416) 3492 2.58 = 0.0215 = 0.022

So, the C.I. is= ( 0.416-0.022, 0.416+0.022) = (0.394, 0.438)

The correct option is B

With 99% confidence, it can be said that the population proportion of adults who say they have started paying bills online last year is between the endpoints of the confidence interval.